OpenStudy (angela210793):

y=1-(sin3x-cos3x)^2 equals with y=sin6x (X belongs to R)....but how do u get it? O.o May some1 pls help me?

OpenStudy (anonymous):

essentially both are the same equation.

OpenStudy (anonymous):

open the squared term and u'll end up getting 2sin3xcos3x=sin6x

OpenStudy (angela210793):

yea i know tht...but how do u transform the 1st into the second?

OpenStudy (angela210793):

really?

OpenStudy (anonymous):

ok,wait

OpenStudy (angela210793):

I got it!!! Yuupiiii!! :) Thnx :)

OpenStudy (anonymous):

1-(sin3x-cos3x)^2=1-(sin^2(3x)+cos^2(3x)-2sin3xcos3x)=1-1+2sin3xcos3x=sin6x . fine ?

OpenStudy (angela210793):

Yea thnx :)